# SN 2024aecx: A double-peaked rapidly evolving Type IIb supernova at 11 Mpc

### v1

@@ 2.2 Spectroscopy
Low-resolution optical spectroscopic observations of SN 2024aecx were initiated with the Himalayan Faint Object Spectrograph Camera (HFOSC) mounted on the Cassegrain focus of 2-meter Himalayan Chandra Telescope (HCT) of the Indian Astronomical Observatory, Hanle, India (Prabhu, 2014). Two grisms Gr#7 (3500–7800 Å) and Gr#8 (5200–9250 Å), available with HFOSC instrument, were used to obtain the spectra, and observations at 12 epochs were obtained between 2024 December 17 (JD 2460662.37) and 2025 February 07 (JD 2460714.41).
@@
We also obtained two epochs of low-resolution optical spectroscopy with the Yunnan Faint Object Spectrograph and Camera (YFOSC) mounted on the Lijiang 2.4-meter Telescope (Wang et al., 2019b). This facility is located near the Mephisto telescope at Lijiang Observatory and is operated by Yunnan Observatories, Chinese Academy of Sciences (CAS). The grism Gr#3, which covers the wavelength range of 3400–9100 Å was used to observe SN 2024aecx on 2024 December 17 (JD 2460662.42) and 2025 December 26 (JD 2460671.36).
The spectra of the arc lamp and spectrophotometric standards were obtained, along with the SN for calibration purposes. The spectroscopic data reduction was carried out in a standard way using IRAF555IRAF stands for Image Reduction and Analysis Facility. It is distributed by the National Optical Astronomy Observatory, which is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation.. After correcting the observed frames for bias and flat fields, one-dimensional spectra were extracted using the optimal extraction method. Wavelength calibration was done by applying the dispersion solutions obtained using arc lamp spectra. The night-sky emission lines were utilized to secure the wavelength calibration. The spectrophotometric standard observed on the same night was used to get the instrumental response curves for flux calibration of the spectra. For those nights where standard star observations were not possible, the response curves obtained during nearby nights were used. The HCT spectra in both grisms (Gr#7, Gr#8) were combined to construct a single flux-calibrated spectrum. The spectra were then scaled with respect to the calibrated magnitudes from Mephisto to bring them to an absolute flux scale. Finally, the SN spectra were corrected for the host galaxy redshift of = 0.002665 0.000002 (from NED). The log of our spectroscopic observations is provided in Table 2.
@@ 3 Results
In this section, we present the detailed light curve properties, the extinction in the SN direction, the color and pseudobolometric bolometric evolution, and the modeling of the light curves.
@@ 3.1 Evolution of light curves
Precursor eruptions are common in interacting events such as SNe IIn and Ibn (Foley et al., 2007; Pastorello et al., 2007; Ofek et al., 2014; Reguitti et al., 2024; Bilinski et al., 2015; Dong et al., 2024b). These eruptions may arise months to years before the SN explosion. However, in a sample of 27 SNe IIb from Palomar Transient Factory, only a marginal detection was observed (Strotjohann et al., 2015). As NGC 3521 was followed up with the Mephisto facility for more than a year, forced photometry was performed at the SN location in multiband images to check any precursor activity before the SN explosion. The limiting magnitudes in different bands are shown in Fig. 2 (left panel). We did not notice a reliable precursor activity before the SN explosion.
@@
The temporal evolution of multi-band light curves is displayed in Figure 2 (right panel). It is evident that there is a decline in all bands just after the explosion, followed by a secondary peak, a typical evolution property of the light curve of SNe IIb (see, for example Richmond et al., 1994; Kumar et al., 2013; Arcavi et al., 2011; Morales-Garoffolo et al., 2014; Tartaglia et al., 2017b; Pellegrino et al., 2023). Although the Swift UVW1, UVM2 band observations are not available after +3 days, a sharp decay can also be seen in those bands. Taking into account densely sampled data in different bands, we used the Gaussian Process Regression method to estimate various parameters of the light curve (tm, tp, mmin, mmax, , ) of SN 2024aecx. Here, tm and tp are the time since the explosion of the light curve minimum after the first peak and the secondary peak, respectively. The mmin and mmax are the magnitudes at tm and tp. The parameters and are the decay in magnitudes in 7 and 15 days after the secondary peak. The data points were sometimes extrapolated in the analysis. These parameters are listed in Table 3. It is inferred that the duration of the shock cooling phase (tm) is longer in the bluer bands than in the redder bands: 7.3 d (-band) and 3.78 d (-band). This is a usual property of double-peaked SNe IIb and has also been seen in SNe 1993J, 2011fu, and 2013df; see Table 4 in Kumar et al. (2013) and Table 6 in Morales-Garoffolo et al. (2014). However, the rise time duration to the secondary peak is opposite, and the bluer bands peaked earlier: 16.72 d (-band) and 22.56 d (-band). Further, it is worth noting that tm (light curve minimum after explosion) of SN 2024aecx is significantly smaller (see Table 3) than SNe 1993J, 2011fu (see Kumar et al., 2013; Morales-Garoffolo et al., 2015) and 2013df (Morales-Garoffolo et al., 2014).
@@
The light curve decay rate in post-peak (second) can be determined with and parameters. A faster decline rate signifies less massive ejecta (shorter diffusion timescales). The values for SN 2024aecx are 2.35 0.05, 1.6 0.5, 1.16 0.04, 1.00 0.05, 1.14 0.04 and 1.40 0.02 mag in bands, respectively (see Table 3). Similarly, the values are also estimated and listed in Table 3. In Fig. 5 (right panel), the and absolute / magnitudes of SE-SNe are plotted. The sample was collected from Taddia et al. (2018b) and the literature. The distribution is scattered and consistent with other studies (Drout et al., 2011; Taddia et al., 2015b, 2018c; Kumar et al., 2022; Gangopadhyay et al., 2023). It is worth noting that SN 2024aecx is a fast-evolving event in the post-peak phase with a decay rate of 0.1 mag day-1 (20–80 days).
@@ 3.2 Extinction towards SN
The reddening due to the Galactic dust along SN 2024aecx line of sight is = 0.05 0.001 mag (Schlafly & Finkbeiner, 2011). Here, the adopted is 3.1 following the extinction law of Cardelli et al. (1989). The fact that SN is located deep inside the host galaxy, strong reddening is expected. Owing to the absence of any reliable technique for estimation, it is challenging to determine the extinction in such a scenario accurately. However, the equivalent width (EW) of Na i D lines can be treated as a good tracer of reddening within the host galaxy (Turatto et al., 2003; Poznanski et al., 2012; Stritzinger et al., 2018) but see Phillips et al. (2013). We estimate the Na i D line EW 2.7 0.5 Å in the spectra near maximum light (–2 d) and applied the relation from Stritzinger et al. (2018): (host) = 0.78 (0.15) EW D [Å] and thus = 0.65 0.10 mag for the host. To further constrain the host galaxy extinction, we tried to match near the maximum light spectra of iPTF13bvn (Cao et al., 2013; Srivastav et al., 2014) with our similar epoch spectra. SN 2024aecx spectra match quite well for value between 0.35 and 0.45 mag (total extinction). Furthermore, for a total = 0.45 mag, the , color evolution of SN 2024aecx matches reasonably well with the average color evolution of SE-SNe templates from the Carnegie Supernova Project compiled by Stritzinger et al. (2018). Therefore, in this study, we consider a total (host and Milky Way) = 0.45 mag, and = 0.65 mag as an upper limit.
3.3 Color indices and Pseudobolometric bolometric light curve
@@
The temperature evolution of SN can be very fast at early times (c.f. adiabatic cooling dominating the LC Woosley et al., 1994) and thus influence different colors. Good cadence data obtained from Mephisto observations were used to estimate the color evolution of SN 2024aecx. The intrinsic colors (corrected for total extinction) and are shown in Fig. 3. The color evolution can be separated into four phases: 0–8, 8–18, 18–40, and beyond 40 days after the explosion. It is evident that all colors indicate an evolution trend of blue-red up to 8 days. Interestingly, SNe 1993J, 2011dh, 2013df, 2016gkg and 2024uwq have also shown a similar color evolution in the bluer bands () at initial epochs (Kumar et al., 2013; Morales-Garoffolo et al., 2014, 2015), although for a shorter duration, 5 days (see Tartaglia et al., 2017b).
@@
The red-blue evolution pattern of SN 2024aecx noticed between days 8 and 18 is rare and only observed in SN 2016gkg and SN 2024uwq. There is a reversal of color () evolution between 18–40 days. Later (after 40 days), the colors and became redder, but no significant variation was observed in color in that phase. The color evolution of SN 2024aecx shows gradual reddening between days 20 to 70 and then follows a blueward trend. It is noted that in our color estimates, the accuracy is 3% after applying the propagation of the error of the photometric measurements. Subrayan et al. (2025) have recently investigated the observational properties of SN 2024uwq (IIb). There is a good resemblance in the evolution of different colors between SN 2024aecx and SN 2024uwq (see their Figure 3).
@@
To construct the pseudobolometric bolometric light curve of SN 2024aecx, we used the bolometric module of the open-source package lightcurve-fitting (Hosseinzadeh et al., 2024). Here, we emphasize that although shorter wavelength magnitudes are available only for a limited duration but ultraviolet radiation significantly contributes to the light curves in the early phases. Therefore, all available magnitudes, including , were utilized for pseudobolometric bolometric light curve estimation. The total extinction towards SN 2024aecx and adopted distance (11.40 Mpc) were used in pseudobolometric bolometric calculations (see Section 3.2).
@@
The result is plotted in Fig. 4 (bottom panel). The initial point of the shock cooling phase has log = 42.60 0.03 erg s-1 and it decreased to a minimum of log = 41.65 0.02 erg s-1 within 5 days after explosion. The second peak was reached afterward on MJD 60679.7 0.3 with log = 42.39 0.02 erg s-1. The blackbody (BB) fitted temperature and radius parameters are also shown in Fig. 4 (top panel). This indicates an abrupt decline in the temperature from 14 kK to 6.5 kK in 8 days, which is consistent with blue colors shown in Fig. 3 in the same duration. Then it gradually reached the secondary maximum (7 kK) and finally settled around 5 kK in the later phases. The shock cooling phase BB temperature evolution of SN 2024aecx resembles nicely with SN 2016gkg (Tartaglia et al., 2017b). There is a gradual change in BB radius, which attains a maximum 12000 around day 30 and slowly declines at later epochs.
In Fig. 5, we compared the absolute band magnitudes of Type Ib and IIb SNe with SN 2024aecx. For the comparison, the data were obtained from the literature. SNe Type Ib used in our comparison are SN 1999dn (Benetti et al., 2011), iPTF13bvn (Srivastav et al., 2014), SN 2005bf (Folatelli et al., 2006; Anupama et al., 2005), SN 2007Y (Stritzinger et al., 2009), SN 2009jf (Valenti et al., 2011), SN 2012au (Pandey et al., 2021) and SN 2017iro (Kumar et al., 2022). Similarly, the compared SNe Type IIb are SN 1993J (Richmond et al., 1994), SN 2008ax (Pastorello et al., 2008), SN 2011dh (Arcavi et al., 2011), SN 2011fu (Kumar et al., 2013; Morales-Garoffolo et al., 2015), SN 2013df (Morales-Garoffolo et al., 2014), SN 2015as (Gangopadhyay et al., 2018), SN 2020acat (Medler et al., 2022; Ergon et al., 2024), SN 2022crv (Gangopadhyay et al., 2023; Dong et al., 2024a), and SN 2022hnt (Farah et al., 2025). The estimated absolute band magnitude for SN 2024aecx at initial epoch and secondary maximum is –17.60 0.10 and –17.94 0.10 mag, respectively. The comparison plot infers that SN 2024aecx is a luminous SN among the SNe IIb sample, and also comparable to the brighter Ib events.

### v2

@@ 2.1 Photometry
Photometry: Aperture photometry was performed on the difference images using SExtractor (E. Bertin & S. Arnouts, 1996a) to construct the light curves. Photometric calibration was carried out using the zero-point values derived from the reference image.
Additionally, we retrieved photometric data ( –bands) from the Asteroid Terrestrial-impact Last Alert System (ATLAS, J. L. Tonry et al., 2018; K. W. Smith et al., 2020). For SN 2024aecx, we downloaded these data from the ATLAS forced photometry service222https://fallingstar-data.com/forcedphot/.
SN 2024aecx was also monitored with Swift/UVOT (N. Gehrels et al., 2004; P. W. A. Roming et al., 2005) in and bands. These observations began on 2024 December 16 (UT) and continued up to 2025 January 13 (UT), providing good coverage. We obtained the publicly available data from the Swift Archives333https://www.swift.ac.uk/swift_portal/ under the ObsID 00018989. As the SN was deep inside the host, it required the exclusion of the contribution from the host galaxy. Fortunately, Swift had previously monitored the same field under ObsID 00084364. We downloaded this data and used it for the host templates. We utilized Swift_host_subtraction444https://github.com/gterreran/Swift_host_subtraction code (P. J. Brown et al., 2009, 2014) to remove the host contribution and perform aperture photometry on the Swift/UVOT images.
The photometric magnitudes in different bands (without extinction correction) are listed in Table 1, and presented in Section 3.1.
(The data used to create this figure are available in the online article.)
@@ 2.2 Spectroscopy
Low-resolution optical spectroscopic observations of SN 2024aecx were initiated with the Himalayan Faint Object Spectrograph Camera (HFOSC) mounted on the Cassegrain focus of 2-meter Himalayan Chandra Telescope (HCT) of the Indian Astronomical Observatory, Hanle, India (T. P. Prabhu, 2014). Two grisms Gr#7 (3500–7800 Å) and Gr#8 (5200–9250 Å), available with the HFOSC instrument (spectral resolution, R 600), were used to obtain the spectra. Spectroscopic observations at 12 epochs were obtained between 2024 December 17 (JD 2460662.37) and 2025 February 07 (JD 2460714.41).
@@
We also obtained two epochs of low-resolution optical spectroscopy with the Yunnan Faint Object Spectrograph and Camera (YFOSC) mounted on the Lijiang 2.4-meter Telescope (C.-J. Wang et al., 2019). This facility is located near the Mephisto telescope at Lijiang Observatory and is operated by Yunnan Observatories, Chinese Academy of Sciences (CAS). The grism Gr#3, which covers the wavelength range of 3400–9100 Å, and has spectral resolution of 300, was used to observe SN 2024aecx on 2024 December 17 (JD 2460662.42) and 2024 December 26 (JD 2460671.36).
The spectra of the arc lamp and spectrophotometric standards were obtained, along with the SN for calibration purposes. The spectroscopic data reduction was carried out in a standard way using IRAF555IRAF stands for Image Reduction and Analysis Facility. It is distributed by the National Optical Astronomy Observatory, which is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation.. The HCT spectra in both grisms (Gr#7, Gr#8) were combined to construct a single flux-calibrated spectrum. The spectra were then scaled with respect to the calibrated magnitudes obtained from Mephisto to bring them to an absolute flux scale. Finally, the SN spectra were corrected for the host galaxy redshift of = 0.002665 0.000002 (from NED). The log of our spectroscopic observations is provided in Table 2. The data are presented in Section 3.4 and included in the electronic version.
@@ 3 Results
In this section, we present the detailed light curve properties, the extinction in the SN direction, the color and pseudobolometric evolution, and the modeling of the light curves.
@@ 3.1 Evolution of light curves
Precursor eruptions are common in interacting events such as SNe IIn and Ibn (R. J. Foley et al., 2007; A. Pastorello et al., 2007; E. O. Ofek et al., 2014; A. Reguitti et al., 2024; C. Bilinski et al., 2015; Y. Dong et al., 2024b). These eruptions may arise months to years before the SN explosion. However, in a sample of 27 SNe IIb from Palomar Transient Factory, only a marginal detection was observed (N. L. Strotjohann et al., 2015). As NGC 3521 was followed up with the Mephisto facility for more than a year, forced photometry was performed at the SN location in multiband images to check any precursor activity before the SN explosion. The limiting magnitudes in different bands are shown in Fig. 2 (left panel). We did not notice reliable flux enhancement in different bands at the SN location during the pre-SN monitoring of the host.
@@
The temporal evolution of multi-band light curves is displayed in Fig. 2 (middle and right panels). It is evident that there is a decline in all bands just after the explosion, followed by a secondary peak, a typical evolution property of the light curve of SNe IIb (see, for example M. W. Richmond et al., 1994; B. Kumar et al., 2013; I. Arcavi et al., 2011; A. Morales-Garoffolo et al., 2014; L. Tartaglia et al., 2017; C. Pellegrino et al., 2023). Although the Swift UVW1, UVM2 band observations are not available after +3 days, a sharp decay can also be seen in those bands (Fig. 2, right panel). Taking into account densely sampled data in different bands, we used the Gaussian Process Regression method to estimate various parameters of the light curve (tm, tp, mmin, mmax, , ) of SN 2024aecx. Here, tm and tp are the time since the explosion of the light curve minimum after the first peak and the secondary peak, respectively. The mmin and mmax are the magnitudes at tm and tp. The parameters and are the decay in magnitudes in 7 and 15 days after the secondary peak. The data points were sometimes extrapolated in the analysis. These parameters are listed in Table 3. It is inferred that the duration of the shock cooling phase (tm) is longer in the bluer bands than in the redder bands: 7.3 d (-band) and 3.78 d (-band). This is a usual property of double-peaked SNe IIb and has also been seen in SNe 1993J, 2011fu, and 2013df; see Table 4 in B. Kumar et al. (2013) and Table 6 in A. Morales-Garoffolo et al. (2014). However, the rise time duration to the secondary peak is opposite, and the bluer bands peaked earlier: 16.72 d (-band) and 22.56 d (-band). Further, it is worth noting that tm (light curve minimum after explosion) of SN 2024aecx is significantly smaller (see Table 3) than SNe 1993J, 2011fu (see B. Kumar et al., 2013; A. Morales-Garoffolo et al., 2015) and 2013df (A. Morales-Garoffolo et al., 2014).
@@
The light curve decay rate in post-peak (second) can be determined with and parameters. A faster decline rate signifies less massive ejecta (shorter diffusion timescales). The values for SN 2024aecx are 2.35 0.05, 1.6 0.5, 1.16 0.04, 1.00 0.05, 1.14 0.04 and 1.40 0.02 mag in bands, respectively (see Table 3). Similarly, the values are also estimated and listed in Table 3. In recent sample based studies, was estimated by F. Taddia et al. (2018b): 1.10 0.10 mag (IIb), 1.25 0.19 mag (Ib), and 1.11 0.45 mag (Ic). Similarly, Ó. Rodríguez et al. (2023) evaluated as 0.88 0.16 mag (IIb), 0.81 0.24 mag (Ib), and 0.86 0.27 mag (Ic). The larger values of SN 2024aecx in different bands indicate that it is one of the fast-evolving events in the post-peak phase than other well-studied SE-SNe.
In Fig. 3, we have compared the absolute and band magnitudes (indicated as in the plot) of Type Ib, Ic, and IIb SNe with SN 2024aecx. The data were obtained from the literature, and the comparison was made with reference to their date of explosion. SNe Type Ib and Ic used in our comparison are SN 1994I (M. W. Richmond et al., 1996), SN 1999dn (S. Benetti et al., 2011), SN 2005bf (G. Folatelli et al., 2006; G. C. Anupama et al., 2005), SN 2007Y (M. Stritzinger et al., 2009), SN 2007gr (D. J. Hunter et al., 2009), SN 2009jf (S. Valenti et al., 2011), SN 2012au (S. B. Pandey et al., 2021), iPTF13bvn (S. Srivastav et al., 2014), SN 2016P (A. Gangopadhyay et al., 2020), SN 2016coi (B. Kumar et al., 2018), SN 2017ein (D. Xiang et al., 2019), SN 2017iro (B. Kumar et al., 2022), and SN 2020oi (J. Rho et al., 2021). Similarly, the compared SNe Type IIb are SN 1993J (M. W. Richmond et al., 1994), SN 2008ax (A. Pastorello et al., 2008), SN 2011dh (I. Arcavi et al., 2011), SN 2011fu (B. Kumar et al., 2013; A. Morales-Garoffolo et al., 2015), SN 2013df (A. Morales-Garoffolo et al., 2014), SN 2015as (A. Gangopadhyay et al., 2018), SN 2020acat (K. Medler et al., 2022a; M. Ergon et al., 2024), SN 2022crv (A. Gangopadhyay et al., 2023; Y. Dong et al., 2024a), and SN 2022hnt (J. R. Farah et al., 2025). The estimated absolute band magnitude for SN 2024aecx at initial epoch and secondary maximum is –17.60 0.10 and –17.94 0.10 mag, respectively. The comparison plot infers that SN 2024aecx is a luminous SN among the SNe IIb sample and is also comparable to the brighter Ib and Ic events.
In Fig. 3 (bottom-right panel), the and absolute / magnitudes of SE-SNe are plotted. The sample was collected from F. Taddia et al. (2018c) and the literature. The distribution is scattered and consistent with other studies (M. R. Drout et al., 2011; F. Taddia et al., 2015b, 2018b; B. Kumar et al., 2022; A. Gangopadhyay et al., 2023). However, SN 2024aecx occupies the extreme right position along with SN 2020oi and SN 1994I in the plot. During the post-peak phase (20–80 days), SN 2024aecx decline with a rate of 0.1 mag day-1, indicating a faster evolution.
@@ 3.2 Extinction towards SN
The reddening due to the Galactic dust along SN 2024aecx line of sight is = 0.05 0.001 mag (Schlafly & Finkbeiner, 2011). Here, the adopted is 3.1 following the extinction law of J. A. Cardelli et al. (1989). The fact that SN is located deep inside the host galaxy, strong reddening is expected. Owing to the absence of any reliable technique for estimation, it is challenging to determine the extinction in such a scenario accurately. To constrain the host galaxy extinction, we use the equivalent width (EW) of the Na i D lines as a proxy for reddening (M. Turatto et al., 2003; M. D. Stritzinger et al., 2018; Ó. Rodríguez et al., 2023). We estimate the Na i D line EW 2.7 0.5 Å in the spectra near maximum light (–2 d) and applied the relation from M. D. Stritzinger et al. (2018): (host) = 0.78 (0.15) EW D [Å] and thus = 0.68 0.22 mag for the host (for = 3.1). To further constrain the host galaxy extinction, we attempted to match the near-maximum light spectra of iPTF13bvn (Y. Cao et al., 2013; S. Srivastav et al., 2014) with our similar epoch spectra. SN 2024aecx spectra match quite well for value between 0.35 and 0.45 mag (total extinction). Additionally, for a total = 0.45 mag, the , color evolution of SN 2024aecx matches reasonably well with the average color evolution of SE-SNe templates from the Carnegie Supernova Project compiled by M. D. Stritzinger et al. (2018). Therefore, we consider a total (host and Milky Way) = 0.45 mag in our analysis, whereas = 0.68 mag as an upper limit.
3.3 Color indices and Pseudo-bolometric light curve
@@
The temperature evolution of SN can be very fast at early times (c.f. adiabatic cooling dominating the LC S. E. Woosley et al., 1994) and thus influence different colors. Good cadence data obtained from Mephisto observations were used to estimate the color evolution of SN 2024aecx. The intrinsic colors (corrected for total extinction) and are shown in Fig. 4. The color evolution can be separated into four phases: 0–8, 8–18, 18–40, and beyond 40 days after the explosion. It is evident that all colors indicate an evolution trend of blue-red up to 8 days. Interestingly, SNe 1993J, 2011dh, 2013df, 2016gkg and 2024uwq have also shown a similar color evolution in the bluer bands () at initial epochs (B. Kumar et al., 2013; A. Morales-Garoffolo et al., 2014, 2015), although for a shorter duration, 5 days (see L. Tartaglia et al., 2017).
@@
The red-blue evolution pattern of SN 2024aecx noticed between days 8 and 18 is rare and only observed in SN 2016gkg and SN 2024uwq. There is a reversal of color () evolution between 18–40 days. Later (after 40 days), the colors and became redder, but no significant variation was observed in the color in that phase. The color evolution of SN 2024aecx shows gradual reddening between days 20 to 70 and then follows a blueward trend. It is noted that in our color estimates, the accuracy is 3 after applying the propagation of the error of the photometric measurements. B. M. Subrayan et al. (2025) have recently investigated the observational properties of SN 2024uwq (IIb). There is a good resemblance in the evolution of different colors between SN 2024aecx and SN 2024uwq (see their Figure 3).
@@
To construct the pseudobolometric light curve of SN 2024aecx, we used the bolometric module of the open-source package lightcurve-fitting (G. Hosseinzadeh et al., 2024). Here, we emphasize that although shorter wavelength magnitudes are available only for a limited duration but ultraviolet radiation significantly contributes to the light curves in the early phases. Therefore, all available magnitudes, including , were utilized for pseudobolometric light curve estimation. The total extinction towards SN 2024aecx and adopted distance (11.40 Mpc) were used in pseudobolometric calculations (see Section 3.2). Here, it is worth mentioning that the bolometric luminosity in SE-SNe is mainly dominated by optical emission, but the UV flux may contribute significantly in early phases. For example, in SN 2024iss the UV emission contributed up to 40 between 0 – 7 days post explosion and decreased rapidly below 10 in a few weeks (M. Yamanaka et al., 2025). Similarly, the NIR emission fraction also varies from 20 (early phases) to 35–40 (1 month after explosion) (K. Medler et al., 2022b; M. Yamanaka et al., 2025). However, for estimating the pseudobolometric luminosity of SN 2024aecx, we only use the UV and optical band fluxes, and the NIR contribution was not included due to the lack of observations in these bands. Therefore, we caution the reader to consider it as a lower limit only.
@@
The result is plotted in Fig. 5 (bottom panel). The initial point of the shock cooling phase has log = 42.60 0.03 erg s-1 and it decreased to a minimum of log = 41.65 0.02 erg s-1 within 5 days after explosion. The second peak was reached afterward on MJD 60679.7 0.3 with log = 42.39 0.02 erg s-1. The blackbody (BB) fitted temperature and radius parameters are also shown in Fig. 5 (top panel). This indicates an abrupt decline in the temperature from 14 kK to 6.5 kK in 8 days, which is consistent with blue colors shown in Fig. 4 in the same duration. Then it gradually reached the secondary maximum (7 kK) and finally settled around 5 kK in the later phases. The shock cooling phase BB temperature evolution of SN 2024aecx resembles nicely with SN 2016gkg (L. Tartaglia et al., 2017). There is a gradual change in BB radius, which attains a maximum 12000 around day 30 and slowly declines at later epochs.